### v1

# Stable Distributed Online Feedback Optimization for Distribution System Voltage Regulation

Abstract
We investigate the distributed voltage regulation problem in distribution systems employing online feedback optimization and short-range communication between physical neighbours. We show that a two-metric approach can be unstable. As a remedy, we propose a nested feedback optimization strategy. Simulation results reveal that while the two-metric approach fails to regulate voltages, the proposed approach achieves even less voltage limit violations than its centralized counterpart.
Index Terms:
Distributed communication, online feedback optimization, voltage regulation, reactive power control
I Introduction
With the proliferation of distributed energy resources (DERs) in distribution systems, voltage regulation has become a challenging task for distribution system operators (DSOs). While local control cannot guarantee the desired voltage regulation [1], model-based feedforward optimization requires accurate grid models, real-time load measurements, and lacks robustness to system uncertainties and disturbances [2].
Recently, online feedback optimization (OFO) has emerged as a promising approach for real-time distribution system control. OFO utilizes measurements as feedback and employs optimization algorithms as feedback controllers to steer the system towards an optimal operating point [2]. While many OFO controllers rely on centralized communication [3, 4], distributed OFO, which relies on communication between physical neighbours, appears to be a compelling approach due to its robustness to single-point failures [5, 1, 6, 7, 8].
While [5, 1, 6] demand particular objective function designs, [7] relies on disseminating global information through neighboring nodes. However, the systematic delays inherent in this information propagation may raise stability issues, potentially undermining its efficacy in practical distribution systems. In this letter, we propose a distributed OFO design that is closely related to that in [8], where the distributed communication is achieved by scaling the gradient with a sparse positive definite matrix $\mathbf{X}^{-1}$. A two-metric approach [9] was further used in [8], where an Euclidean projection step was used to ensure satisfaction of the DER capacity constraint. The approach is called two-metric because the matrix $\mathbf{X}^{-1}$ is used for the gradient scaling while the identity matrix $\mathbf{I}$ is used for the projection.
However, utilizing such a two-metric approach might lead to divergence, as shown in [9]. Instead of the Euclidean distance, the distance measured in terms of the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ should be used in the projection step to ensure convergence [9]. Solving this non-Euclidean projection problem would however require global information, contradicting the distributed OFO design. To tackle this challenge, we propose an iterative algorithm that requires only local information and operates in an online feedback-based manner. This leads to our nested feedback optimization approach, where the outer loop comprises the OFO iterations yielding tentative but not necessarily feasible DER setpoints, and the inner loop solves the non-Euclidean projection to map the tentative setpoints to actual feasible setpoints.
II Online feedback optimization
II-A Problem formulation
Consider the distribution system voltage regulation using reactive power as in Problem (1). DERs are located in the set $\mathcal{N}$, and $q_{i}$ represents their reactive power injections. The objective function (1a) minimizes the reactive power cost, where $c_{i}$ is the cost parameter. Constraint (1b) enforces the lower and upper voltage limits $\underline{v}$ and $\overline{v}$, where $\mu_{i}$ and $\lambda_{i}$ are the respective dual variables. Finally, (1c) ensures that the reactive power injection $q_{i}$ respects the lower and upper limits $\underline{q}_{i}$ and $\overline{q}_{i}$. Note that the voltage $v_{i}$ is dependent on $q_{i}$ and other disturbances in the system such as active power injections of the DERs and active and reactive power consumption of non-controllable load units.
$\underset{q_{i},\forall i\in\mathcal{N}}{\operatorname{minimize}}$ $\quad f:=\frac{1}{2}c_{i}q_{i}^{2}$ (1a)
$\operatorname{s.t.}$ $\quad\underline{v}\leq v_{i}\leq\overline{v}:\mu_{i},\lambda_{i},\forall i\in\mathcal{N}$ (1b)
$\underline{q}_{i}\leq q_{i}\leq\overline{q}_{i},\forall i\in\mathcal{N}$ (1c)
II-B Centralized primal-dual gradient projection algorithm
Solving Problem (1) offline in a feedforward way demands an accurate grid model and disturbance data to evaluate $v_{i}$, which might not be available. The feedforward approach also lacks robustness [2]. To tackle these challenges, OFO has emerged as a promising approach for real-time distribution system management. This section introduces an instance of OFO based on the primal-dual gradient projection (PDGP) algorithm [3], which requires a centralized gather-and-broadcast communication architecture. The PDGP algorithm performs projected gradient ascent and descent steps iteratively for the dual and primal variables, respectively. At each time step (iteration) $k$, it repeats the following three steps:
  1. 1.
    For each bus $i\in\mathcal{N}$, collect its voltage measurement $\tilde{v}_{i}^{k}$ and update $\lambda_{i}$ and $\mu_{i}$, where the projection operator is defined as $[u]^{+}:=\max(u,0)$ and $\alpha$ with various superscripts represents step sizes:
    $\lambda_{i}^{k+1}\leftarrow\big{[}\lambda_{i}^{k}+\alpha^{\lambda}(\tilde{v}_{i}^{k}-\overline{v})\big{]}^{+}$ (2a)
    $\mu_{i}^{k+1}\leftarrow\big{[}\mu_{i}^{k}+\alpha^{\mu}(\underline{v}-\tilde{v}_{i}^{k})\big{]}^{+}$ (2b)
  2. 2.
    The DSO broadcasts the updated dual variables $\boldsymbol{\lambda}^{k+1}$ and $\boldsymbol{\mu}^{k+1}$ to all buses, where $\boldsymbol{\lambda}^{k+1}$ and $\boldsymbol{\mu}^{k+1}$ are vectors constructed by stacking $\lambda_{i}^{k+1}$ and $\mu_{i}^{k+1}$, respectively.
  3. 3.
    For each bus $i\in\mathcal{N}$, update its reactive power setpoints $q_{i}$, where $[u]_{\underline{q}_{i}}^{\overline{q}_{i}}$ represents the Euclidean projection into $[\underline{q}_{i},\overline{q}_{i}]$, and $\mathbf{X}$ is a $|\mathcal{N}|$-dimensional dense matrix mapping reactive power injections to voltages, i.e., the sensitivity matrix, as constructed in [8]. Because of the non-sparsity of $\mathbf{X}$, Step 2 is required to provide information needed in Step 3.
    $q_{i}^{k+1}\leftarrow\big{[}q_{i}^{k}-\alpha[c_{i}q_{i}^{k}+[\mathbf{X}(\boldsymbol{\lambda}^{k+1}-\boldsymbol{\mu}^{k+1})]_{i}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}$ (2c)
II-C Two-metric distributed approach
To avoid a centralized communication architecture and promote the distributed communication architecture where communication is only between physical neighbours in the system, [8] proposed to scale the gradient in (2c) by a sparse positive definite matrix $\mathbf{X}^{-1}$, which has non-zero entries only in the diagonal and the $i$-th row $j$-th column if nodes $i$ and $j$ are adjacent. This yields (3) and replaces (2c) in the OFO, where $\mathbf{q}^{k}$ is the stack of $q^{k}_{i}$, and $\mathbf{C}$ is a $|\mathcal{N}|$-dimensional diagonal matrix with elements $c_{i}$ defined in (1a).
$q_{i}^{k+1}\leftarrow\big{[}q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}$ (3)
Computing (3) for node $i$ now requires only information from its physcial neighbours. We stack $q_{i}^{k}-\alpha[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}$ for all nodes (tentative but not necessarily feasible setpoints to be projected) as the vector $\mathbf{\dot{q}}^{k+1}$ and define $\mathcal{X}^{k}$ as $[\underline{q}_{1}^{k},\overline{q}_{1}^{k}]\times[\underline{q}_{2}^{k},\overline{q}_{2}^{k}]\times\cdots\times[\underline{q}_{|\mathcal{N}|}^{k},\overline{q}_{|\mathcal{N}|}^{k}]$ for further developments, where $\times$ is the Cartesian product of sets.
II-D Proposed distributed approach
The two-metric approach in Section II-C does not necessarily lead to algorithm convergence [9]. To ensure convergence, the projection in (3) should use the distance measured in terms of the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ due to the use of the scaling matrix $\mathbf{X}^{-1}$ [9], that is:
$\mathbf{q}^{k+1}\leftarrow\underset{\mathbf{u}\in\mathcal{X}^{k}}{\operatorname{argmin}}\quad\frac{1}{2}(\mathbf{u}-\mathbf{\dot{q}}^{k+1})^{\intercal}\mathbf{X}(\mathbf{u}-\mathbf{\dot{q}}^{k+1})$ (4)
This however contradicts the distributed communication architecture of the overall OFO algorithm since directly solving (4) again requires global information. To tackle this challenge, this letter proposes an iterative algorithm that requires only local information and is implemented in an online feedback-based manner. This leads to our nested feedback optimization approach, where the outer loop comprises the OFO iterations giving tentative reactive power setpoints, and the inner loop solves the non-Euclidean projection problem (4). Our key observation is that given a feasible $\mathbf{u}^{\tau}$, the gradient of the objective function in (4) is $\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})$, which can be approximated by the voltage difference between $\mathbf{v}(\mathbf{u}^{\tau})$ and $\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ assuming system disturbances do not change significantly when all nodes are locally implementing this inner iterative algorithm. The algorithm is based on the non-scaled Euclidean gradient projection and proceeds as:
  1. 1.
    Implement $\mathbf{q}^{k}_{\epsilon}:=\mathbf{q}^{k}+\epsilon(\mathbf{\dot{q}}^{k+1}-\mathbf{q}^{k})$ and measure the voltage $\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$. Estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$ as $\mathbf{\tilde{v}}^{k}$ + ($\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$-$\mathbf{\tilde{v}}^{k}$)/$\epsilon$.
  2. 2.
    Given a feasible $\mathbf{u}^{\tau}$, implement it, measure the voltage $\mathbf{\tilde{v}}(\mathbf{u}^{\tau})$, compute the next implementable reactive power setpoint $\mathbf{u}^{\tau+1}$ as (5), and repeat this step for $T$ iterations.
    $\mathbf{u}^{\tau+1}\leftarrow[\mathbf{u}^{\tau}-\alpha^{u}(\mathbf{\tilde{v}}(\mathbf{u}^{\tau})-\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1}))]_{\mathbf{\underline{q}}^{k}}^{\mathbf{\overline{q}}^{k}}$ (5)
Note that these two steps can be implemented for each node using only local information. In the first step, $\epsilon$ represents a small exploration parameter, which is similar to that in model-free feedback optimization [10]. Since $\mathbf{\dot{q}}^{k+1}$ can be pretty far away from the feasible set, it might be impossible to implement it. Therefore, we implement $\mathbf{q}^{k}_{\epsilon}$ and estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$ based on it. This could still result in a minor temporary capacity limit violation, which should be manageable by the DER. To ensure strict capacity limit satisfaction, the original feasible set can be slightly deflated [10]. The initial feasible $\mathbf{u}^{\tau}$ in the second step can be simply chosen as $\mathbf{q}^{k}$.
III Simulation results
In our simulation experiment, we compare the centralized OFO based on the gather-and-broadcast communication, the distributed two-metric approach in [8], and our proposed nested distributed approach in both static and dynamic cases. We study a 100% penetration scenario of photovoltaics (PVs) in a 96-node low-voltage distribution system, which experiences overvoltage issues. Details of the distribution system and the generation and load data with a 6-second resolution can be found in [11] and references therein. The PV inverters are assumed to be oversized by 20% to provide sufficient reactive power. The reactive power cost matrix $\mathbf{C}$ is chosen as the identity matrix. The lower and upper voltage limits are 0.95 and 1.05 pu, respectively. The unit for power is $kW/kVar$. The step sizes are chosen with a trial-and-error strategy [8] as $\alpha^{\lambda}=\alpha^{\mu}=10^{6}$, $\alpha=10^{-1}$ (for centralized), $\alpha=10^{-3}$ (for distributed), and $\alpha^{u}=10^{2}$. The exploration parameter $\epsilon$ is chosen as $10^{-5}$, resulting in a maximum DER reactive power limit violation of 2%. For each generation and load data point, an OFO iteration is run which yields tentative reactive power setpoints, followed by one exploration step and 5 inner feedback iterations of (5) ($T=5$).
Figure 1: Voltage profiles over 200 iterations for two remote buses in the static case using the proposed and the two-metric approaches. We count both the outer and inner iterations for our proposed approach.
Figure 2: Reactive power profiles after 200 iterations for the centralized, the proposed, and the two-metric approaches.
Figure 3: Voltage profiles for the most remote bus in the dynamic simulation. The average voltage violation (AVV) values are $1.9\times 10^{-4}$, $1.3\times 10^{-4}$, and $4.6\times 10^{-3}$ pu for the centralized, the proposed, and the two-metric approaches, respectively, which is defined as $\operatorname{AVV}=\frac{1}{K}\sum_{k=1}^{K}[v^{k}-\overline{v}]^{+}$, where $k$ is the iteration index and $K$ is the number of iterations.
In the static case, the generation and load data are kept unchanged to study the convergence of various approaches. Figure 1 shows that while the two-metric approach does not converge and fails to regulate voltages, our proposed approach converges quickly after approximately 154 iterations, which correspond to around 22 outer and 132 inner iterations. We also compare the final reached reactive power setpoints after 200 iterations. Figure 2 shows that the centralized and our proposed approaches converge to the same solution. This is reasonable since the gradient scaling only alters the trajectory, but does not change the solution to Problem (1).
In the dynamic case, we study the performance of the approaches in a time-varying environment, where the approaches are implemented online without reaching convergence. Figure 3 shows the significant voltage limit violation without control and demonstrates that our proposed nested distributed approach successfully enforces the voltage limit even for the most remote, and therefore, most sensitive bus, achieving an average voltage violation (AVV) of $1.3\times 10^{-4}$ pu, which is better even than that of the centralized approach ($1.9\times 10^{-4}$ pu). We conjecture that this is because of the smoothing brought by the inner feedback loop. The two-metric approach, however, fails to regulate voltages, resulting in 34 times higher AVV.
IV Conclusion
In this letter, we focused on the distributed voltage regulation problem in distribution systems. We showed that the previously proposed two-metric approach might fail. As a remedy, we proposed a nested feedback optimization approach. Our simulation results showed that the approach achieved even better voltage regulation than a centralized approach while only requiring short-range communication between physical neighbours. The proposed distributed approach however does not extend to congestion management while the centralized does so. The approach works for either active or reactive power control, but not joint control. This is because the observation in Section II-D does not hold anymore since the voltage difference is no longer the gradient of either active or reactive power injection, but represents their joint non-separable impact. However, one might deploy active and reactive power feedback controllers separately at different sampling time. Whether the system under two feedback controllers is stable is left for future investigations.