### v3

# Distributed Online Feedback Optimization for Real-time Distribution System Voltage Regulation

Abstract
We investigate the real-time voltage regulation problem in distribution systems employing online feedback optimization (OFO) with short-range communication between physical neighbours. OFO does not need an accurate grid model nor estimated consumption of non-controllable loads, affords fast calculations, and demonstrates robustness to uncertainties and disturbances, which render it particularly suitable for real-time distribution system applications. However, many OFO controllers require centralized communication, making them susceptible to single-point failures. This paper proposes a distributed OFO design based on a nested feedback optimization strategy and analyzes its convergence. Numerical study results demonstrate that the proposed design achieves effective voltage regulation and outperforms other distributed and local approaches.
Index Terms:
Distributed communication, online feedback optimization, reactive power control, voltage regulation
I Introduction
The rapid growth of distributed energy resources (DERs) has introduced significant uncertainty and volatility to distribution systems. With the high resistance/reactance ratios of distribution cables, voltage fluctuations are occurring more frequently. Consequently, real-time voltage regulation has become a significant operational challenge for distribution system operators (DSOs). The conventional approach of DSOs is grid reinforcement, but this demands significant investments, a skilled workforce, and considerable time. During this lengthy process, a complementary solution is to exploit the flexibility of network assets and DERs. In this context, online feedback optimization (OFO) has recently emerged as a promising strategy, e.g. in [1, 2, 3, 4, 5, 6].
OFO utilizes measurements as feedback and employs optimization algorithms as feedback controllers to steer physical systems towards their optimal operating points [7], which are defined by the well-established optimal power flow (OPF) problem. Compared to directly solving the OPF problem in a feedforward manner, OFO does not need an accurate grid model nor data of non-controllable loads, affords fast calculations, and is robust against uncertainties and disturbances in distribution systems due to its feedback-based nature [8, 7, 9]. Figure 1 illustrates feedforward and feedback systems.
Figure 1: Block diagrams of feedforward and feedback systems for voltage regulation using reactive power, where $q$ represents system input (reactive power), $v$ represents system output (voltages), $\omega$ ($\widehat{\omega}$) represents (estimated) system disturbance (non-controllable loads), and $h$ maps system input and disturbance to system output. This figure is adapted from [10].
OFO also brings distinct advantages over other existing approaches. Compared to local droop control [11, 3, 12], OFO ensures more reliable voltage regulation [3] and can pursue grid-level objectives. In contrast to deep reinforcement learning [13, 14], OFO does not need a complicated (often centralized offline) training process while offering theoretical guarantees to obtain globally optimal solutions. Furthermore, unlike using traditional voltage regulators such as on-load tap changers and capacitor banks [15], OFO provides rapid voltage regulation with inverter-interfaced DERs and can complement these slow-reacting devices in real time [16].
While many OFO controllers rely on centralized communication [1, 17, 2, 8, 4, 5, 18, 9, 19, 20, 21], distributed OFO based on communication between physical neighbours appears to be a compelling approach due to its robustness to single-point failures[6, 3, 16, 22, 23, 24]. Among the existing distributed designs, [6, 3, 16] demand particular forms of objective functions, while [22] relies on disseminating global information through neighbouring nodes. However, the systematic delays inherent in this information propagation may raise stability issues, potentially undermining its efficacy in practical distribution systems. In [23], the distributed design was achieved by only keeping components in the network sensitivity matrix related to local and neighbouring nodes. This heuristic approach does not assure adequate voltage regulation.
In this paper, we propose a distributed OFO design inspired by [24], where the distributed implementation was enabled by the sparsity of the inverse of the network sensitivity matrix $\mathbf{X}$, i.e. $\mathbf{X}^{-1}$ is sparse. Both our proposed and their approaches continue with scaling the gradient by $\mathbf{X}^{-1}$ followed by a projection step to ensure satisfaction of the DER capacity constraint. It is conceptually attractive to use the Euclidean distance in the projection, which can be implemented using only local information. However, as shown in [25], the distance measured by the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ should be used instead in the projection step to ensure descent iterations and thus convergence. With the Euclidean distance, it would be a two-metric approach which can lead to algorithm divergence [25]. Solving this non-Euclidean projection problem would however require global information since $\mathbf{X}$ is dense, contradicting the distributed OFO design. As a remedy, [24] introduced additional dualization of the DER capacity constraint on top of the Euclidean projection to ensure the DER constraint satisfaction. Nevertheless, this strategy takes a prohibitively large number of iterations to converge, which may jeopardize its performance in online implementation.
We tackle this challenge by proposing an iterative algorithm to solve this non-Euclidean projection problem that requires only local information and operates in an online feedback-based manner. This leads to our overall nested feedback optimization approach, where the outer loop comprises the OFO iterations requiring communication between physical neighbours and yielding tentative but not necessarily feasible DER setpoints, while the inner loop solves the non-Euclidean projection problem to map the tentative setpoints to actual feasible setpoints using local information. To summarize, the main contributions of this paper are:
  • A local iterative algorithm is proposed to solve the non-Euclidean projection problem that operates in an online feedback-based manner.
  • An efficient nested feedback optimization strategy that ensures adequate voltage regulation while eliminating the need for an accurate grid model and consumption of non-controllable loads is developed.
  • The proposed strategy is theoretically analyzed and numerically assessed on a 96-bus system with high-resolution generation and load data.
The remainder of this paper is structured as follows: Section II presents the system modeling, problem formulation, and a centralized OFO strategy. Section III introduces the proposed nested approach and analyzes its convergence. Section IV presents a numerical study, while Section V draws conclusions and discusses future work.
II Problem formulation and centralized OFO
II-A Distribution system modeling
[…]
The linearized DistFlow equations in (1) were proposed in [26] and can be used to model balanced radial distribution systems. Equations (1a)-(1b) represent nodal active and reactive power balance constraints, respectively. The voltage relation is modeled in (1c), which further leverages the assumption that $v_{i}^{2}-v_{j}^{2}\approx 2(v_{i}-v_{j})$ since $v_{i}\approx 1\text{ pu},\forall i\in\mathcal{N}$.
$P_{ij}+p_{j}=\sum_{k:(j,k)\in\mathcal{E}}P_{jk}+p^{d}_{j},\forall j\in\mathcal{N}^{+},$ (1a)
$Q_{ij}+q_{j}=\sum_{k:(j,k)\in\mathcal{E}}Q_{jk}+q^{d}_{j},\forall j\in\mathcal{N}^{+},$ (1b)
$v_{i}-v_{j}=r_{ij}P_{ij}+x_{ij}Q_{ij},\forall(i,j)\in\mathcal{E}.$ (1c)
Applying (1) directly in a feedforward voltage regulation scheme requires estimates of $p_{j}^{d}$ and $q^{d}_{j}$, $\forall j\in\mathcal{N}^{+}$. Moreover, the modeling inaccuracy will also result in voltage approximation errors. To synthesize the proposed feedback controller for voltage regulation, we leverage (1) to derive network sensitivities, which relate changes in nodal active and reactive power injections to voltage variations. Following the steps in [27] or [28], the linear relation in (2) can be constructed.
$\mathbf{v}=v_{0}\mathbf{1}+\mathbf{R}(\mathbf{p}-\mathbf{p}^{d})+\mathbf{X}(\mathbf{q}-\mathbf{q}^{d}).$ (2)
The $N\times N$-dimensional symmetrical matrices $\mathbf{R}$ and $\mathbf{X}$ are given in (3), where $\mathcal{E}_{i}$ represents the set of cables on the unique path from the substation bus to bus $i$.
$\mathbf{R}_{ij}=\sum_{(h,k)\in\mathcal{E}_{i}\cap\mathcal{E}_{j}}r_{hk},\mathbf{X}_{ij}=\sum_{(h,k)\in\mathcal{E}_{i}\cap\mathcal{E}_{j}}x_{hk}.$ (3)
These matrices accordingly capture the network sensitivities, i.e.
$\frac{\partial\mathbf{v}}{\partial\mathbf{p}}=\mathbf{R},\frac{\partial\mathbf{v}}{\partial\mathbf{q}}=\mathbf{X}.$ (4)
When the resistances and reactances of the cables are all positive, [27] shows that $\mathbf{R}$ and $\mathbf{X}$ are both symmetrical positive definite. Furthermore, [24] shows that their inverse matrices $\mathbf{R}^{-1}$ and $\mathbf{X}^{-1}$ have the sparsity pattern such that
$\mathbf{R}_{ij}^{-1}\neq 0,\mathbf{X}_{ij}^{-1}\neq 0\iff(i,j)\in\mathcal{E}\text{ or }i=j.$ (5)
Figure 2 visualizes $\mathbf{X}$ and $\mathbf{X}^{-1}$ for the test system in Section IV, where the sparsity patterns are shown.
(a) $\mathbf{X}$ (pu/kVar)
(b) $\mathbf{X}^{-1}$ (kVar/pu)
Figure 2: Heat maps of the network sensitivity matrix $\mathbf{X}$ and its inverse $\mathbf{X}^{-1}$.
II-B Problem formulation
Consider the problem of distribution system voltage regulation using reactive power as formulated in (6). The objective function (6a) minimizes the reactive power use cost, where $c_{i}>0$ is the cost parameter. Constraint (6b) enforces the lower and upper voltage limits $\underline{v}$ and $\overline{v}$, where $\mu_{i}$ and $\lambda_{i}$ are the respective dual variables. Finally, (6c) ensures that the reactive power injection $q_{i}$ respects the lower and upper limits $\underline{q}_{i}$ and $\overline{q}_{i}$, which capture DER capacity and power factor constraints. It is noteworthy that $v_{i}$ is dependent on the decision variable $q_{i}$ and other disturbances in the system such as active and reactive power consumption of non-controllable loads.
$\underset{q_{i},\forall i\in\mathcal{N}^{+}}{\operatorname{minimize}}$ $\quad\sum_{i\in\mathcal{N}^{+}}\frac{1}{2}c_{i}q_{i}^{2},$ (6a)
$\operatorname{s.t.}$ $\quad\underline{v}\leq v_{i}\leq\overline{v}:\mu_{i},\lambda_{i},\forall i\in\mathcal{N}^{+},$ (6b)
$\underline{q}_{i}\leq q_{i}\leq\overline{q}_{i},\forall i\in\mathcal{N}^{+}.$ (6c)
For further developments, the dual problem of (6) is formulated in (7).
$\underset{\mu_{i}\geq 0,\lambda_{i}\geq 0,\forall i\in\mathcal{N}^{+}}{\operatorname{maximize}}\big{\{}\underset{\underline{q}_{i}\leq q_{i}\leq\overline{q}_{i},,\forall i\in\mathcal{N}^{+}}{\operatorname{minimize}}\quad\mathcal{L}(\mathbf{q},\boldsymbol{\mu},\boldsymbol{\lambda})\big{\}},$ (7)
where the partial Lagrangian function is defined in (8).
$\mathcal{L}(\mathbf{q},\boldsymbol{\mu},\boldsymbol{\lambda})=\sum_{i\in\mathcal{N}^{+}}\left[\frac{1}{2}c_{i}q_{i}^{2}+\lambda_{i}\left(v_{i}-\overline{v}\right)+\mu_{i}\left(\underline{v}-v_{i}\right)\right].$ (8)
II-C Centralized primal-dual gradient projection algorithm
Solving (6) offline in a feedforward way demands an accurate grid model and disturbance data to evaluate $v_{i}$, which might not be readily available. The feedforward approach also lacks robustness [7]. Thus, this section introduces a feedback controller based on OFO. The controller leverages primal-dual gradient projection (PDGP) as the underlying algorithm [1, 17, 2] to track the time-varying optimizers of (6), which requires a centralized gather-and-broadcast communication architecture. Specifically, the PDGP algorithm performs projected gradient ascent and descent steps iteratively for the dual and primal variables, respectively. At each time step (iteration) $k$, it repeats the following three steps:
  1. 1.
    For each bus $i\in\mathcal{N}^{+}$, collect its voltage measurement $\tilde{v}_{i}^{k}$ and update $\lambda_{i}$ and $\mu_{i}$ using (9a)-(9b), where the projection operator is defined as $[u]^{+}=\max(u,0)$, $\alpha$ with various superscripts represents step sizes, while $r^{d}$ and $r^{p}$ in (9c) are the dual and primal regularization factors [1, 17, 2], respectively.
    $\lambda_{i}^{k+1}=\big{[}\lambda_{i}^{k}+\alpha^{d}(\tilde{v}_{i}^{k}-\overline{v}-r^{d}\lambda_{i}^{k})\big{]}^{+},$ (9a)
    $\mu_{i}^{k+1}=\big{[}\mu_{i}^{k}+\alpha^{d}(\underline{v}-\tilde{v}_{i}^{k}-r^{d}\mu_{i}^{k})\big{]}^{+}.$ (9b)
  2. 2.
    The DSO broadcasts the updated dual variables $\boldsymbol{\lambda}^{k+1}$ and $\boldsymbol{\mu}^{k+1}$ to all buses.
  3. 3.
    For each bus $i\in\mathcal{N}^{+}$, update its reactive power setpoints $q_{i}$ using (9c), where $[u]_{\underline{u}}^{\overline{u}}$ represents the Euclidean projection into $[\underline{u},\overline{u}]$. Since $\mathbf{X}$ is dense, Step 2 is needed to provide information for Step 3.
    $q_{i}^{k+1}=\big{[}q_{i}^{k}-\alpha[c_{i}q_{i}^{k}+[\mathbf{X}(\boldsymbol{\lambda}^{k+1}-\boldsymbol{\mu}^{k+1}+r^{p}\mathbf{q}^{k})]_{i}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}.$ (9c)
III Distributed OFO
III-A Two-metric distributed approach
To avoid the centralized communication requirement and enable the distributed communication architecture where communication is only between physical neighbours, [24] proposed to scale the gradient in (9c) by the sparse positive definite matrix $\mathbf{X}^{-1}$, which has non-zero entries only in the diagonal and the $i$-th row $j$-th column if buses $i$ and $j$ are adjacent, i.e. as in (5). Continuing with the Euclidean projection yields (10) which replaces (9c) in the OFO, where $\mathbf{C}$ is a $N$-dimensional diagonal matrix with $c_{i}$ defined in (6a).
$q_{i}^{k+1}=\big{[}q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}+r^{p}q_{i}^{k}]\big{]}_{\underline{q}_{i}^{k}}^{\overline{q}_{i}^{k}}.$ (10)
Computing (10) for node $i$ now requires only information exchange with its physical neighbours. However, the two-metric approach does not necessarily lead to algorithm convergence since it is not in general a descent iteration [25]. This is illustrated in Fig. 3 with an example. For further developments, we define
$\mathbf{\dot{q}}^{k+1}=$ (11)
$\quad\left[q_{i}^{k}-\alpha[[\mathbf{X}^{-1}\mathbf{C}\mathbf{q}^{k}]_{i}+\lambda_{i}^{k+1}-\mu_{i}^{k+1}+r^{p}q_{i}^{k}],i\in\mathcal{N}^{+}\right],$
which are tentative but not necessarily feasible setpoints to be projected for all nodes, and define
$\mathcal{X}^{k}=[\underline{q}_{1}^{k},\overline{q}_{1}^{k}]\times[\underline{q}_{2}^{k},\overline{q}_{2}^{k}]\times\cdots\times[\underline{q}_{N}^{k},\overline{q}_{N}^{k}],$ (12)
where $\times$ is the Cartesian product of sets.
Figure 3: An illustration showing that the two-metric approach is not a descent iteration, where $\mathbf{x}^{k}$ is the current iterate which is also the optimum, $X$ is the feasible region, and $\nabla f^{k}$ is the gradient at $\mathbf{x}^{k}$. With (plain) gradient projection on the left panel, the iterate stays at the optimum, i.e. $\mathbf{x}^{k+1}=\mathbf{x}^{k}$. With the two-metric approach on the right panel where $\mathbf{Y}$ represents the gradient scaling metric, $\mathbf{x}^{k+1}$ leaves the optimum.
III-B Proposed distributed approach
It has been shown in [25] that, to ensure convergence, the projection should use the distance measured in terms of the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ due to the scaling matrix $\mathbf{X}^{-1}$ [25], that is:
$\mathbf{q}^{k+1}=\underset{\mathbf{u}\in\mathcal{X}^{k}}{\operatorname{argmin}}\quad\frac{1}{2}(\mathbf{u}-\mathbf{\dot{q}}^{k+1})^{\intercal}\mathbf{X}(\mathbf{u}-\mathbf{\dot{q}}^{k+1}).$ (13)
This however contradicts the distributed communication architecture of the overall OFO algorithm since, unlike the Euclidean projection, solving (13) again requires global information. As a solution, [24] proposed dualization of the DER capacity limits, combined with the Euclidean projection, to ensure the DER constraint satisfaction. However, this approach requires an excessively large number of iterations to reach convergence, significantly hindering its performance in online implementation.
Figure 4: Illustration of the approximation.
To tackle this challenge, this paper proposes an iterative algorithm to solve (13) that requires only local information and is implemented in an online feedback-based manner. This leads to our nested feedback optimization approach, where the outer loop comprises the OFO iterations giving tentative reactive power setpoints, and the inner loop solves the non-Euclidean projection problem (13) to generate feasible setpoints. Our key observation is that given a feasible $\mathbf{u}^{\tau}\in\mathcal{X}^{k}$, the gradient of the objective function in (13) at $\mathbf{u}^{\tau}$ is $\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})$, which can be approximated by the voltage difference $\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ according to (2) assuming system disturbances do not change significantly when all nodes are locally implementing this inner iterative algorithm, i.e.
$\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})\approx\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1}).$ (14)
The proposed algorithm is based on the non-scaled Euclidean gradient projection and proceeds as:
  1. 1.
    Implement
    $\mathbf{q}^{k}_{\epsilon}=\mathbf{q}^{k}+\epsilon(\mathbf{\dot{q}}^{k+1}-\mathbf{q}^{k}),$ (15)
    and measure the voltage $\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$. Estimate
    $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})\approx\mathbf{\tilde{v}}^{k}+(\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})-\mathbf{\tilde{v}}^{k})/\epsilon.$ (16)
  2. 2.
    Given a feasible $\mathbf{u}^{\tau}$, implement it, measure the voltage $\mathbf{\tilde{v}}(\mathbf{u}^{\tau})$, compute the next implementable reactive power setpoint $\mathbf{u}^{\tau+1}$ as (17), and repeat this step for $T$ iterations.
    $\mathbf{u}^{\tau+1}=[\mathbf{u}^{\tau}-\alpha^{u}(\mathbf{\tilde{v}}(\mathbf{u}^{\tau})-\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1}))]_{\mathbf{\underline{q}}^{k}}^{\mathbf{\overline{q}}^{k}}.$ (17)
Remark 1.
In the first step, $\epsilon$ represents a small exploration parameter, which is similar to that in model-free feedback optimization [29, 30, 31, 32]. Since $\mathbf{\dot{q}}^{k+1}$ can be pretty far away from the feasible set, it might be impossible to implement it. Therefore, we implement $\mathbf{q}^{k}_{\epsilon}$ and estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$ based on it. This is illustrated in Fig. 4. This could still result in minor temporary DER capacity limit violations, which should be manageable by the DERs. To ensure strict capacity limit satisfaction, the original feasible set can be slightly deflated [32]. The initial feasible $\mathbf{u}^{\tau}$ in the second step can be chosen as $\mathbf{q}^{k}$. The online implementation of the overall nested algorithm is summarized in Algorithm 1.
Data: $\mathbf{q}^{0}$
Result: $\mathbf{q}^{k}$, $k\geq 0$; $\mathbf{u}^{\tau}$, $\tau\leq T$
1 Set k = 0;
2 while $k\geq 0$ do
3       Implement $\mathbf{q}^{k}$ and collect voltage measurement $\tilde{\mathbf{v}}(\mathbf{q}^{k})$;
4       Communicate $\mathbf{q}^{k}$ between physical neighbours and calculate $\dot{\mathbf{q}}^{k+1}$ via (11);
5       Implement $\mathbf{q}^{k}_{\epsilon}$, measure voltage $\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$ and estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$;
6       Set $\tau=0$;
7       while $\tau<T$ do
8             Implement $\mathbf{u}^{\tau}$ and measure voltage $\tilde{\mathbf{v}}(\mathbf{u}^{\tau})$ (choose $\mathbf{u}^{0}$ as $\mathbf{q}^{k}$);
9             Calculate next implementable $\mathbf{u}^{\tau+1}$ via (17);
10             Set $\tau=\tau+1$;
11
12       end while
13      Set $\mathbf{q}^{k+1}=\mathbf{u}^{T}$;
14       Set $k=k+1$.
15 end while
Algorithm 1 Nested feedback optimization algorithm
III-C Convergence analysis
The convergence analysis of the nested algorithm is based on the following two assumptions.
Assumption 1.
The linear relation (2) holds between voltages and reactive power injections.
Based on the first assumption, $\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ is the exact gradient of the objective function in (13) at $\mathbf{u}^{\tau}$. The inner loop then represents a standard Euclidean gradient projection algorithm to solve the quadratic program (13). Consequently, $\mathbf{u}^{\tau}$} is a converging sequence by invoking Proposition 2.3.2 in [25] with $0<\alpha^{u}<2/\lambda_{max}(\mathbf{X})$ where $\lambda_{max}(\mathbf{X})$ is the largest eigenvalue of $\mathbf{X}$.
Assumption 2.
The inner iterations reach convergence after $T$ iterations.
The overall nested algorithm thus represents a scaled PDGP algorithm to solve (6). By a transformation of variables defined by: $\mathbf{q}^{\prime}=\mathbf{X}^{\frac{1}{2}}\mathbf{q}$ and $\mathbf{v}^{\prime}=\gamma\mathbf{v}$ where $\gamma=\sqrt{\frac{\alpha^{d}}{\alpha}}$ is the squared root of the ratio between the dual and primal step sizes, the problem (6) can be cast as an equivalent quadratic program in the space of $\mathbf{q}^{\prime}$ and $\mathbf{v}^{\prime}$ as in (18).
$\underset{\mathbf{q^{\prime}}}{\operatorname{minimize}}$ $\quad\frac{1}{2}\mathbf{q^{\prime}}^{\intercal}\mathbf{X}^{-\frac{1}{2}}\mathbf{C}\mathbf{X}^{-\frac{1}{2}}\mathbf{q^{\prime}},$ (18a)
$\operatorname{s.t.}$ $\quad\gamma\mathbf{\underline{v}}\leq\mathbf{v^{\prime}}\leq\gamma\mathbf{\overline{v}}:\boldsymbol{\mu^{\prime}},\boldsymbol{\lambda^{\prime}},$ (18b)
$\mathbf{\underline{q}}\leq\mathbf{X}^{-\frac{1}{2}}\mathbf{q^{\prime}}\leq\mathbf{\overline{q}}.$ (18c)
The unscaled regularized PDGP algorithm to solve (18) includes the following iterations:
$\boldsymbol{\lambda^{\prime}}^{k+1}=[\boldsymbol{\lambda^{\prime}}^{k}+\alpha(\gamma\mathbf{\tilde{v}}-\gamma\mathbf{\overline{v}}-r^{d}\boldsymbol{\lambda^{\prime}}^{k})]^{+},$ (19a)
$\boldsymbol{\mu^{\prime}}^{k+1}=[\boldsymbol{\mu^{\prime}}^{k}+\alpha(\gamma\mathbf{\underline{v}}-\gamma\mathbf{\tilde{v}}-r^{d}\boldsymbol{\mu^{\prime}}^{k})]^{+},$ (19b)
$\mathbf{q^{\prime}}^{k+1}=\underset{\mathbf{q^{\prime},\mathbf{\underline{q}}\leq\mathbf{X}^{-\frac{1}{2}}\mathbf{q^{\prime}}\leq\mathbf{\overline{q}}}}{\operatorname{argmin}}\lVert\mathbf{q^{\prime}}-\big{[}\mathbf{q^{\prime}}^{k}-\alpha[\mathbf{X}^{-\frac{1}{2}}$ (19c)
$\qquad\cdot\mathbf{C}\mathbf{X}^{-\frac{1}{2}}\mathbf{q^{\prime}}^{k}+\gamma\mathbf{X}^{\frac{1}{2}}(\boldsymbol{\lambda^{\prime}}^{k+1}-\boldsymbol{\mu^{\prime}}^{k+1})+r^{p}\mathbf{q^{\prime}}^{k}]\big{]}\rVert_{2}^{2}.$
which are equivalent to (9a)-(9b) and (11)-(13) [25, Eqs. (2.37) and (2.38)]. Invoking Theorem 3 in [2] with $\alpha$ chosen using [2, Eq. (34)], (19) then converges to the optimizer of the regularized saddle-point problem of (18). Consequently, our proposed nested algorithm also converges to the optimizer of the regularized saddle-point problem of (6).
Remark 2.
First, while the linear voltage-reactive power relation is assumed to synthesize the controller and analyze its convergence, AC power flow is run in the simulation to calculate the actual voltages. Such an assumption was also made in the proofs of [3, 23]. Second, a large $T$ is not necessary in the simulation. We use $T=4$ so the inner loop might not converge. However, the algorithm performs well in the simulation. Notably, a similar assumption was made in [3].
IV Case study
IV-A Case description
Figure 5: A 96-bus test system from Simbench [33].
In the numerical simulation, we study a 100% penetration scenario of photovoltaics (PVs) in a 96-bus system, i.e. we assume every node has a PV installation. The test system is adopted from Simbench [33] and is shown in Fig. 5, which is expected to experience overvoltage issues. The network topology and impedance data are the same as those in the original dataset. For each bus, a PV installation with a randomly generated DC capacity from 3-10 kW is assumed. The PV inverters are assumed to be oversized by 20% to provide sufficient reactive power. For the dynamic case study, generation and load data with a 6-second resolution are used. The PV profiles are linearly interpolated using the HelioClim-3 dataset [34] with a 1-minute resolution. Base load profiles are obtained by aggregating 1-second resolution data from the ECO dataset [35]. AC power flow problems are solved with a high-performance library PowerGridModel [36].
The reactive power cost matrix $\mathbf{C}$ is chosen as the identity matrix. The lower and upper voltage limits are 0.95 and 1.05 pu, respectively. The unit for power is kW/kVar. The step sizes are chosen with a trial-and-error strategy [24] as $\alpha^{d}=10^{6}$, $\alpha=5\times 10^{-4}$, and $\alpha^{u}=10^{2}$. The exploration parameter $\epsilon$ is chosen as $10^{-5}$, resulting in a maximum DER reactive power limit violation of 0.6%. The primal and dual regularization is used to establish the theoretical convergence result. The impact of some small regularization parameters on numerical results is negligible. In the simulation, we set them as $10^{-4}$. For each generation and load data point, an OFO iteration is run which yields tentative reactive power setpoints, followed by one exploration step and 4 inner feedback iterations of (17), i.e. we assume that every setpoint in Algorithm 1 is implemented for 1 second. Finally, the nested algorithm requires only basic arithmetic operations and is computationally very efficient. For a series implementation of a 4-hour simulation, i.e. in total 14400 iterations, it takes 8.1 seconds, averaging 0.56 milliseconds per iteration.
Figure 6: Voltage profiles over 500 iterations for several buses in the static case using the centralized and the proposed approaches. We count both the outer and inner iterations for our proposed approach.
Figure 7: Reactive power profiles after 500 iterations for the centralized and the proposed approaches.
(a) No control
(b) Centralized
(c) Proposed
(d) Two metric
(e) Tang 2024 [23]
(f) Qu 2020 [24]
(g) Droop control [37, 38]
(h) Summary for bus 95
Figure 8: Voltage profiles under different approaches in the dynamic simulation.
(a) Centralized
(b) Proposed
(c) Two metric
(d) Tang 2024 [23]
(e) Qu 2020 [24]
(f) Droop control [37, 38]
Figure 9: Reactive power profiles under different approaches in the dynamic simulation.
IV-B Static case results
In the static case, the generation and load data are kept unchanged to study the convergence of the nested approach. The simulation uses data at 12:00. Figure 6 shows that the proposed approach suppresses the initial voltage limit violations quickly and converges after around 200 total iterations, i.e. 33 outer OFO iterations. The approach also utilizes the available voltage limit as efficiently as the centralized one. Figure 7 compares the converged reactive power setpoints, which shows that the centralized and the proposed approaches converge to the same solution. This is reasonable since the gradient scaling only alters the converging trajectory, but does not change the solution to (6).
IV-C Dynamic case results
In the dynamic case, we study the performance of the approaches in a time-varying environment, where the approaches are implemented online to track the time-varying optimizers of (6) without waiting for them to converge. Figure 8a shows significant voltage limit violations without control. Figures 8b and 8c demonstrate that the centralized and our proposed distributed nested feedback optimization approaches both successfully enforce the voltage limit, although temporary voltage limit violations are inevitable due to the corrective nature of these algorithms. The simple two-metric approach and the heuristic approach in [23] result in prolonged voltage limit violations as seen in Figs. 8d and 8e, respectively. Figure 8f shows that the distributed approach in [24] does not track the time-varying optimizers well and experiences oscillations at the end of the simulation period. The discrepancy between the actual setpoints and algorithm iterates (without projection) potentially renders the algorithm slow and unstable. Figure 8g shows that the local droop control approach does not enforce the voltage limit due to lack of coordination. Finally, for clarity, Fig. 8h presents voltages for the most remote, and therefore, the most sensitive bus (bus 95) under different approaches.
Furthermore, Fig. 9 shows reactive power profiles under different approaches. Shown in Figs. 9a and 9b, our proposed distributed approach achieves similar profiles with the centralized one, with an average setpoint deviation of $-1.3\times 10^{-3}$ kVar. The two-metric approach and the heuristic approach do not work well in this case study. The approach in [24] experiences oscillations. The local droop control approach does not in theory or practice track the optimizers of (6).
Finally, we define the average voltage violation (AVV) in (20), where $k$ is the iteration index and $K$ is the total number of iterations.
$\operatorname{AVV}=\frac{1}{K}\sum_{k=1}^{K}\left([\tilde{v}^{k}-\overline{v}]^{+}+[\underline{v}-\tilde{v}^{k}]^{+}\right).$ (20)
TABLE I: Average voltage violations for various approaches.
Approach AVV (pu) w.r.t. centralized
No control $1.5\times 10^{-2}$ 192x
Centralized $8.0\times 10^{-5}$ 1x
Proposed $1.6\times 10^{-4}$ 2x
Two-metric $2.5\times 10^{-3}$ 31x
Tang 2024 [23] $3.3\times 10^{-3}$ 41x
Qu 2020 [24] $8.6\times 10^{-4}$ 11x
Droop control [37, 38] $4.5\times 10^{-4}$ 6x
Table I shows the AVV values for bus 95 (the most sensitive bus) under various approaches. An AVV of $1.6\times 10^{-4}$ pu is achieved for the proposed approach, which is only higher than that of the centralized approach. The latter however requires a centralized communication architecture and is thus susceptible to single-point failures. The proposed nested feedback approach significantly outperforms other existing distributed or local approaches.
V Conclusion
In this paper, we focused on the distributed voltage regulation problem in distribution systems and proposed a nested feedback optimization approach. We theoretically analyzed its convergence. Our simulation results showed that the approach, while only requiring short-range communication between physical neighbours, achieved satisfactory voltage regulation and outperformed existing distributed and local approaches.
The proposed distributed approach does not extend to congestion management while the centralized does so. The current presentation is based on radial systems but the approach can be readily extended to meshed systems using the approach in [3] to construct the sensitivity matrix $\mathbf{X}$. Extending the approach to unbalanced systems is an interesting future direction, where the focus is to construct a positive definite $\mathbf{X}$ with a sparse inverse matrix. Finally, the approach works for either active or reactive power control, but not joint control. This is because the observation in Section III-B does not hold anymore since the voltage difference is no longer the gradient of either active or reactive power injection, but represents their joint non-separable impact. However, one might deploy active and reactive power feedback controllers separately at different sampling time. Whether the system under two feedback controllers is stable is left for future investigations.